OpenStudy (anonymous):

What are the zeros of the function f(x) = f of x equals three x squared minus nine x plus six divided by quantity three x minus three ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{3x^2 - 9x+6}{3x-3}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

plug in f(x) = 0 and solve for x

OpenStudy (anonymous):

thats it ?

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{3x^2 - 9x+6}{3x-3}\] \[\Large 0 = \frac{3x^2 - 9x+6}{3x-3}\] \[\Large 0(3x-3) = 3x^2 - 9x+6\] \[\Large 0 = 3x^2 - 9x+6\] \[\Large 3x^2 - 9x+6 = 0\] \[\Large 3(x^2 - 3x+2) = 0\] \[\Large 3(x-2)(x-1) = 0\] I'll let you finish up

OpenStudy (anonymous):

is it 2?

jimthompson5910 (jim_thompson5910):

that's one solution there's one more

jimthompson5910 (jim_thompson5910):

3(x-2)(x-1) = 0 turns into x-2 = 0 or x-1 = 0 x = 2 or x = ???

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

correct x = 2, x = 1 are your two roots/zeros

OpenStudy (anonymous):

ok so for my homework they have both 1 and 2 which would be my answer ?

jimthompson5910 (jim_thompson5910):

yes, both are part of the solution set

OpenStudy (anonymous):

but i can only pick one answer

jimthompson5910 (jim_thompson5910):

hmm odd, there's no "select all that apply" ?

OpenStudy (anonymous):

no theres -1, 1, 2, and -2

jimthompson5910 (jim_thompson5910):

oh I'm not thinking lol

jimthompson5910 (jim_thompson5910):

in the denominator, we have 3x-3 if x = 1, then 3x-3 = 3*1-3 = 3-3 = 0 you cannot divide by zero, so x = 1 is NOT in the domain

jimthompson5910 (jim_thompson5910):

only x = 2 is the solution

OpenStudy (anonymous):

oh ok thank you next question !